OpenStudy (anonymous):

x: (4x - 6)(2x - 5) = 0

OpenStudy (anonymous):

You will use the distributive property for this, but taking the 4x times the 2x to make it a 8x^2. Then the -6 times the 2x for a -12x. The 4x time the -5 to make a -20x. Then -6 time the -5 to make positive 30.

OpenStudy (anonymous):

then you still need to simplify the problem

OpenStudy (anonymous):

This will leave you with the answer 8x^2-32x+30

jimthompson5910 (jim_thompson5910):

if you want to solve for x, then use the zero product property (4x - 6)(2x - 5) = 0 4x - 6 = 0 or 2x - 5 = 0 now solve each equation for x

OpenStudy (anonymous):

nevermind, didn't see the zero xD

OpenStudy (anonymous):

deleting

jimthompson5910 (jim_thompson5910):

that's ok, happens to me all the time but you provided a good way to FOIL or expand out

OpenStudy (anonymous):

haha thank you :)

jimthompson5910 (jim_thompson5910):

yw